OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 45 mi/h = ___ ft/s a. 66 c. 26.4 b. 15.6 d. 19.2

OpenStudy (dumbcow):

"dimensional analysis" haha that makes you sound smart....in other words "unit conversion" \[\frac{45 mi}{hr} *\frac{1 hr}{3600 s}*\frac{5280 ft}{mi} = \frac{45*5280}{3600}\frac{ft}{s}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

so wat is the answer

OpenStudy (anonymous):

66